Rose is 4/5 as old as Teena. Six years ago, she was 2/3 as old as Teena. How old are they?
Answer by lwsshak3(11628)

Rose is 4/5 as old as Teena. Six years ago, she was 2/3 as old as Teena. How old are they?
**
let x=Teena's present age
4x/5=Rose's present age
..
6 yrs ago:
x-6=Teena's age
4x/5-6=Rose's age
Rose=2/3 of Teena's age
(4x/5)-6=(2/3)(x-6)
LCD:15
12x-90=10(x-6)
12x-90=10x-60
2x=30
x=15
4x/5=60/5=12
Teena's present age: 15
Rose's present age:12